a) Write an inequality that represents the story problem below.
b) Solve the inequality to find the amount of time you can rent the bicycle. c) Graph your solution on a numberline.
You want to rent a bicycle to use on the Falling Waters Trail. There is an initial $6.00 fee that allows you to use the bike for 3 hours. There is an additional fee of $2.00/hour for every hour after that. You have $15.00 that you can spend on renting the bicycle.

Answer by Theo(13342)

the formula will be:
6 + 2 * (x - 3) <= 15
x represents the number of hours you can rent the bike.
15 represents the amount of money you have available for renting the bike for x hours.
subtract 6 from both sides of this equation to get:
2 * (x - 3) <= 9
simplify to get:
2 * x - 6 <= 9
add 6 to both sides of this equation to get:
2 * x <= 15
divide both sides by 2 to get:
x <= 15/2
simplify to get:
x <= 7.5
the maximum number of hours you can ren the bike is 7.5.
you pay 6 dollars up front.
this gets you the first 3 hours.
you have 9 dollars left.
at 2 dollars an hour, you can get an additional 4.5 hours,.
3 hours plus 4.5 hours = 7.5 hours total for your 15 dollars.
your solution on the number line will start at 0 and end at 7.5, inclusive of both end points.